OpenStudy (anonymous):

Which ratio is equivalent to 3:5? a. 6/10 b. 1/5 c. 10/5 d. 15/5

OpenStudy (alexandervonhumboldt2):

simplify all ratios

OpenStudy (alexandervonhumboldt2):

hint: 6/10=(3*2)/(5*2) 10/5=2 15/5=3

OpenStudy (nerdsarecool):

15/5 right

OpenStudy (alexandervonhumboldt2):

it is wrong

OpenStudy (alexandervonhumboldt2):

15/5=3 how is 3=3:5?